# VISER: Visually-Informed System for Enhanced Robustness in Open-Set Iris Presentation Attack Detection

### 18 Mar 2026 v1

@@ 4.4 Observations
To further contextualize our results, we report on observations from our experiments that may help inform state-of-the-art iris PAD model design. For example, AUROC results in Table 1 show that the D-NetPAD with cross entropy baseline performs quite well on the ”contacts + print” attack type to the point that no other salience-based methods are able to improve upon its already superior performance. This high performance is somewhat explained by the similarity of the ”printout” and ”contacts + print” attack types being very similar. On the other hand, the ”artificial” class shows the most room for improvement as every method but Low Entropy hand annotations surpassed the baseline with all forms of eye tracking showing large performance increases and De-noised Initial ET showing the single largest improvement across the board. Eye tracking methods as a whole showed broad improvement in the categories of ”printout” and ”synthetic” indicating that human feature selection in these categories is making a noticeable improvement.
@@
Finally, we report on the performance of the DINOv2 foundation models. Unlike past experiments that showed superior closed-set performance, the performance of DINOv2 models is mixed. In Table 1, only DINOv2 + SVM-RBF was able to exceed the XENT baseline while the other two variants, SVM-Linear and LogReg, failed to surpass the baseline performance. Furthermore, while SVM-RBF shows solid improvement in ”artificial”, ”synthetic”, and ”post-mortem” attack types, SVM-Linear and LogReg performance is explained by a large improvement in ”artificial” and equivalent or worse than baseline performance in all other categories. In Table 2, we see that all three methods were near equivalent or worse than the XENT baseline with all three methods performing very well on ”post-mortem” samples and generally poorly on all other attack types.
@@ 5 Conclusion
This paper proposes VISER, a methodology for open-set iris PAD detection leveraging different methods and forms of saliency, including two distinct human saliency forms, for saliency-guided training. In our study, we compared the performance of these methods in AUROC and APCER at BPCER 1% against a D-NetPAD cross entropy baseline and reported on the improvement, or absence of, on this baseline. Our findings indicate that the best performing variant was De-noised Initial ET heatmaps with an average AUROC improvement of 0.0608 and an APCER at BPCER 1% improvement of 0.1063 over D-NetPAD indicating there is an noticeable performance improvement from de-noising eye tracking maps prior to saliency-guided training. The remaining eye tracking methods also showed noticeable improvements over D-NetPAD in both metrics whereas the majority of the other methods failed to make an impact. Neither the hand annotations, nor the segmentation masks, were able to improve over D-NetPAD in APCER at BPCER 1% and only the Equal Entropy hand annotations were able to marginally improve AUROC. Finally, we evaluated DINOv2 with three different classification methods inspired by prior work in closed set evaluations. Our results show that while DINOv2 + SVM-RBF showed an AUROC improvement of 0.0485, the other two methods, LogReg and SVM-Linear, failed to improve the baseline AUROC. In APCER at BPCER 1%, only LogReg made a marginal improvement over D-NetPAD while SVM-Linear and SVM-RBF failed to improve. We reported in our observations and limitations that we cannot truly guarantee open-set testing for DINOv2 models since we do not know its training data but that its performance increases largely occur on one attack type. Finally we discuss some limitations and discussed ways for future researchers to adapt this work. To facilitate follow up research efforts, the source codes, model weights, and saliency maps are offered along with this paper.

### 07 May 2026 v2

@@ 4.4 Observations
To further contextualize our results, we report on observations from our experiments that may help inform state-of-the-art iris PAD model design. In Table 1, none of the methods were able to improve APCER at BPCER 1% performance on the ”contacts + print” attack type as the XENT baseline is already a strong performer at 0.02337. While DINOv2-Large struggled across the board, SigLIP2 and DINOv3 configurations showed strong performance particularly on Post-Mortem, Diseased, and Synthetic samples. When viewing specific configurations in table 2, DINOv3 + CNN and SigLIP2 + SVM-RBF showed strong performance with both methods improving the baseline for 5 out of 7 attack types and of those improvements, all were significant. Eye tracking methods also made large improvements with all configurations making significant improvements on at least 3 attack types and no significant performance degradation comparative XENT. In particular, De-noised Initial ET significantly increased performance on 4 out of 7 attack types without significantly degrading performance on the remaining three. Furthermore, De-noised Initial ET recorded the best pooled mean AP/BPCER 1% and the lowest 95% confidence threshold of all tested methods with similar numbers for all eye tracking configurations.
Finally, we report on the performance between eye tracking methods and foundation models. While foundation model configurations, SigLIP2 + SVM-RBF and DINOv3 + CNN, achieve the highest magnitude of improvement ($|d|scores>2.469$) with significant improvement on 5 out of 7 attack types, their performance is volatile across attack types as exhibited by their significant performance degradation on the remaining 2 attack types, printouts and printouts of subjects wearing textured contact lens. Conversely, eye tracking methods recorded the highest pooled AP/BPCER 1% mean values, significantly improved 3-4 attack types, and crucially did not significantly degrade performance on any attack type tested. These findings suggest that eye tracking saliency can effectively regularize the model towards robust features whereas the foundation models tested may overfit to specific high-level features.
4.5 Limitations
We acknowledge some limitations of this work, which we address in this section. Foundation models were tested in our experiments due to their rising popularity and impressive capabilities in the iris PAD domain. However, we cannot guarantee open-set testing for any foundational model, including the ones tested here, as we do not know which iris images, if any, were absorbed into the model training. In terms of performance, our results show that collectively eye tracking saliency maps outperformed hand annotations head-to-head while improving the DenseNet121 cross entropy baseline. Since collecting human saliency is expensive and time consuming, we set out to investigate if one type is better than the other for saliency-guided training. While our results show that eye tracking is the better candidate in open-set iris PAD, we acknowledge that further exploration in other domains is needed to validate a broader claim. Likewise, our results indicate that de-noising eye tracking maps led to an increase in performance. However, this was largely done to mitigate concerns of a continuous noisy signal commonly associated with eye tracking. Future studies should explore de-noising other eye tracking datasets and consider different clustering parameters depending on the task at hand before a broader claim can be made as to whether de-noising eye tracking saliency maps has a significant impact on saliency-guided training compared to standard eye tracking gaze maps.
4.6 Observations
To further contextualize our results, we report on observations from our experiments that may help inform state-of-the-art iris PAD model design. For example, AUROC results in Table LABEL:tab:compact_auroc_grouped show that the D-NetPAD with cross entropy baseline performs quite well on the ”contacts + print” attack type to the point that no other salience-based methods are able to improve upon its already superior performance. This high performance is somewhat explained by the similarity of the ”printout” and ”contacts + print” attack types being very similar. On the other hand, the ”artificial” class shows the most room for improvement as every method but Low Entropy hand annotations surpassed the baseline with all forms of eye tracking showing large performance increases and De-noised Initial ET showing the single largest improvement across the board. Eye tracking methods as a whole showed broad improvement in the categories of ”printout” and ”synthetic” indicating that human feature selection in these categories is making a noticeable improvement.
@@ 4.6 Observations
Finally, we report on the performance of the DINOv2 foundation models. Unlike past experiments that showed superior closed-set performance, the performance of DINOv2 models is mixed. In Table LABEL:tab:compact_auroc_grouped, only DINOv2 + SVM-RBF was able to exceed the XENT baseline while the other two variants, SVM-Linear and LogReg, failed to surpass the baseline performance. Furthermore, while SVM-RBF shows solid improvement in ”artificial”, ”synthetic”, and ”post-mortem” attack types, SVM-Linear and LogReg performance is explained by a large improvement in ”artificial” and equivalent or worse than baseline performance in all other categories. In Table 1, we see that all three methods were near equivalent or worse than the XENT baseline with all three methods performing very well on ”post-mortem” samples and generally poorly on all other attack types.
@@ 5 Conclusion
In this study, we compared the performance of methods and forms of saliency, including two distinct human saliency forms, for saliency-guided training in the iris PAD task. Specifcially, we made comparisons of APCER at BPCER 1% against a DenseNet121 cross entropy baseline and reported on the improvement, or absence of, on this baseline. Our findings indicate that the best performing variant was De-noised Initial ET heatmaps with the best mean pooled AP/BPCER 1% of 0.6765 and significantly improving performance on 4 out of the 7 attack types without significantly decreasing performance on the others. The remaining eye tracking methods also showed noticeable improvements with all variants surpassing the mean pooled AP/BPCER 1% of all other tested configurations including the XENT baseline. Conversely, the majority of the other methods failed to make an impact as neither the hand annotations, nor the segmentation masks, were able to improve over the baseline. Finally, we evaluated three foundation models, DINOv2-Large, DINOv3, and SigLIP2, each with three different classification methods inspired by prior work in closed set evaluations. Our results show that while some configurations of SigLIP2 and DINOv3 made significant improvements over the baseline, they also actively harmed the baseline in other attack types and ultimately did not exceed the performance of the eye tracking-based methods im mean pooled AP/BPCER 1%. We reported in our observations and limitations that we cannot truly guarantee open-set testing for foundation models since we do not know what data they are trained on but that their performance increases largely occur on a few attack types and operate in a ”High- Risk, High-Reward” paradigm. Finally we discuss some limitations and discussed ways for future researchers to adapt this work. To facilitate follow up research efforts, the source codes, model weights, and saliency maps are offered along with this paper.